A father is six times as old
as the son. Four years ago
the father was ten times
as old as his son.
Find the
1. son's age
2. Father's age
3. Son's age in 3yrs times.

Answer by mananth(16946)

son's age =x
father's age =6x
4 years ago
son's age = x-4
father's age =6x-4
conditon
6x-4=10(x-4)
6x-4=10x-40
6x-10x=-40+4
-4x=-36
/-4
x=9
son's age = 9 years
father's age = 6x= 54
son's age 3 years later --> x+3= 9+3=12